OpenStudy (anonymous):

Need help ASAP before 3:45 PLz!!!!!!!!

OpenStudy (anonymous):

In a water treatment plant, water passes through a cone-shaped filter with a height of 4 m and a diameter of 9 m. The water flows from the filter into a cylinder below it. One full cone-shaped filter fills one-fifth of the cylinder. What are the dimensions of the cylinder? Use 3.14 to approximate pi. A. h = 3 m; r = 3 m B. h = 4.23 m; r = 5 m C. h = 15 m; r = 6 m D. h = 33.75 m; r = 2 m I think D

OpenStudy (anonymous):

h = 33.75 m; r = 2 m

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

Thx one more?

OpenStudy (anonymous):

\frac{9/2*9/2*4* \pi }{3} = \frac{81*3.14 }{3} =27*3.14 (the cone's v) so multiply 5 to the v = 135*3.14 (cylinder's v) A:3*3*3*3.14=27*3.14 X B:5*5*4.23*3.14=105.75*3.14 X C:6*6*15*3.14=540*3.14 X D:2*2*33.75*3.14=135*3.14 O

OpenStudy (anonymous):

A cone-shaped container is filled with liquid. The container has a radius of 60 cm and an height of 210 cm. The liquid is drained from the container at a rate of 1099 cm3 per hour. How many hours will it take to drain all of the liquid? Use 3.14 to approximate pi. ___h

OpenStudy (anonymous):

@viacomplys

OpenStudy (anonymous):

720

OpenStudy (anonymous):

Sure?

OpenStudy (anonymous):

yes senpai

OpenStudy (anonymous):

Thx c;

OpenStudy (anonymous):

anytime senpai

OpenStudy (anonymous):

Dont call me that lol can u help with like 5 more?

OpenStudy (anonymous):

I a new thread tho?

OpenStudy (anonymous):

*in